OpenStudy (anonymous):

Given the data set (200, 150, 325, 350, 100, 75), Find the mean of the data set.

OpenStudy (anonymous):

\[\frac{200 + 150 + 325 + 350 + 100 + 75}{6}=\frac{1200}{6}=200\]